Two long parallel wires are 10cm apart and carry currents of 6A and 4 A. What is the field at B at 4A carrying wire due to 6A wire
1
Expert's answer
2020-05-15T08:59:06-0400

magnetic field ="\frac{\mu_oi}{2\pi r}=\frac{4\pi\times10^{-7}\times6}{2\pi\times0.1}=12\times10^{-6}=1.2\times10^{-5}"